OpenStudy (anonymous):

I bought a dozen eggs. the weights in grams are as follow: 54,55,57,59,59,61,63,66,66,67,67,77 grams. what is the probability I randomly pick 2 eggs that weight over 133 grams?

OpenStudy (kropot72):

None of the 12 eggs weigh over 133 grams. Therefore the probability of randomly picking 2 eggs that weigh over 133 grams is zero. Have you typed the question correctly?

OpenStudy (anonymous):

sorry i ment randomly picking two, that added together, weight >133g

OpenStudy (anonymous):

so like 77+67 would be 144g

OpenStudy (campbell_st):

well you need to find the number of possible 2 egg choices 12 x 11 = than the number where the combined weight > 133 the P( combined weight > 133) = number of combined > 133 ------------------------ total number of arrangements

OpenStudy (kropot72):

The order in the selection of each pair of eggs does not affect the combined weight of the 2 eggs. Therefore the number of combinations of the 12 eggs taken 2 at a time is \[12C2=\frac{12!}{2!10!}=\frac{12\times 11}{2}\]

OpenStudy (anonymous):

i agree with that. Im having trouble coming up with a clear way of thinking about the >133 part

OpenStudy (anonymous):

like if 77 then 9 choices if 67 then 4 choices (66,66,67,77) if 66 then 3 choices(67,67,77)

OpenStudy (anonymous):

16/66? im afraid im overcounting some of them

OpenStudy (campbell_st):

yep thats the way to look at it

OpenStudy (anonymous):

no thats wrong because i counted 77 twice

OpenStudy (anonymous):

i think anyway

OpenStudy (kropot72):

My take on combinations weighing more than 133 grams: Egg weight Combinations 54 0 55 0 57 1 59 1 59 1 61 1 63 1 66 1 66 1 67 1 67 1 67 + 67 1

OpenStudy (kropot72):

@sahsah The question states that the combined weight of the 2 eggs is over 133 grams (not equal to 133 grams).